The doctor's order reads: Drug Y is 2g per 1000mL D5W at 4mg per minute. The drop factor is 60gtts per mL. What is the flow rate in drops per minute?
Answer #1

ANSWER.

* Calculation of flow rate in drops per minute.

          Formula

               Flow rate(gtt/min) = Volume(ml) /Time(min) * Drop factor.

        Volume(ml) of drug per minute.

              Formula = Desired dose /Dose in hand *Quantity.

                 Desired dose = 4 mg.

                 Dose in hand = 2g.

        Convert gm to mg for calculation

         So

             Dose in hand = 2000 mg.

                  Quantity = 1000 ml.

          As per the above formula

               Volume(ml) of drug per minute = 4/2000 * 1000

                                                                 = 0.002 * 1000

                                                                  = 2 ml.

     So

        Flow rate (gtt/min) = volume/time * drop factor.

                                       = 2/1 * 60

                                       = 120 gtt/min.